OpenStudy (anonymous):

Solve by completing the square x^2 +5/2 x=5 What is the solution? x=

OpenStudy (anonymous):

\[x^2 +\frac{5}{2} x=5\]

OpenStudy (anonymous):

\[x^2 + 5/2x + 25/16 = 5 + 25/16\] \[(x + 5/4) ^2 = 105/ 16\] \[x+ 5/4 = \pm \sqrt{105}/4\] \[x = (\pm \sqrt{105} - 5) / 4\]